<p>Both GSI vacuum group and edinburgh ion gauges OFF</p>
<p>Fees power cycled. All system checks OK.</p>
<p>Writing to file, alpha background overnight.</p>
<p>Moved detector IN 2mm. Noise appears reduced with any distance moved in so overnight spectra to be taken while IN. Not moving further in as possible beam tomorrow.</p>
<p>22:44</p>
<p>Leakage current stable, slowly decreasing ~3.57 uA (attachment 1)</p>
<p>Statistics look OK</p>
<p>Temperatures OK</p>
<p>System wide check all pass - OK</p>
<p>Noise appears much reduced on spectrum Browser channels. (attachement 2). Compared to same plot on elog 197</p>
<p>XYtotal per pixel (attachment 3)</p>
<p>Ex histogram (attachment 4)</p>
<p> </p> |